# RoboMIND: Benchmark on Multi-embodiment Intelligence Normative Data for Robot Manipulation

### v2

@@ I Introduction
Beyond establishing such a large-scale and diverse dataset, we conduct extensive experiments to not only validate the dataset’s effectiveness but also evaluate various algorithms’ performance, providing a comprehensive benchmark analysis. Specifically, we evaluate the task success rates using single-task imitation learning methods, including ACT [112], Diffusion Policy [14], and BAKU [36]. Additionally, we assess the generalization capabilities and task success rates of Vision-Language-Action (VLA) large models such as OpenVLA [73], RDT-1B [61], and CrossFormer [21]. The experimental results demonstrate that RoboMIND can be effectively utilized by various single-task imitation learning algorithms and successfully adapted to VLA large models. The high-quality information provided by our dataset enables successful task execution across different approaches in real-world scenarios. Furthermore, pre-training the entire VLA models using the full RoboMIND dataset results in significant improvements in task performance across multiple robot types.
@@ III-A Data Collection and Storage
Teleoperation is widely applied in the data collection processes for various types of robots [81, 110, 104, 37, 59, 102, 80, 13, 96]. Different types of robots also have specific teleoperation devices for data collection. For example, researchers typically use VR headsets and motion capture suits to collect humanoid robot motion data. They capture the state of human movements and map this motion onto the humanoid robot platform, enabling the robot to replicate these movements while simultaneously collecting a comprehensive dataset [13, 96]. RoboMIND contains teleoperation data from various types of robots, such as single-arm robots (Franka Emika Panda [31], UR5e [84]), dual-arm robots (AgileX Cobot Magic V2.0 [83]), and humanoid robots (X-Humanoid Tien Kung [9]).
@@
Figure 4: We define 8 quality assurance criteria in the data collection process. Touch Excess: Unnecessary contact with objects by the robotic arm; Movement not Smooth: Noticeable jerking or interruptions in robotic arm movements; Secondary Grabbing: Repeated grasping attempts after failures in robotic arm operations; Mechanical Arm Shaking: Abnormal vibrations in the robotic arm; Collision before Grabbing: Collision of the gripper with surrounding objects before grasping; Image Distortion: Data collection quality issues; Failed Placement: Incorrect placement of objects; Gripper out of the Camera: Frames in which the gripper exceeds video frame boundaries. We show 8 trajectory examples that failed to pass the quality assurance due to different reasons. Each example includes three images that depict the dynamic process of the trajectory. We use red boxes and markers to highlight the reasons for failure.
@@ III-C Data Annotation
While the visual and robot proprioceptive information can be extracted directly from the collected videos and trajectories, we need to provide better semantic information from the data to aid model training. For each collection task, its detailed and accurate linguistic descriptions are provided. These linguistic annotations can be utilized for training currently popular VLA models. RoboMIND collection tasks encompass numerous long horizon tasks, where a uniform linguistic description may be insufficient to capture the full complexity and nuances of the entire task. Thus, we offer detailed fine-grained linguistic annotations for each movement occurring within a trajectory, as illustrated in Figure 5. We annotate 10k successful robot motion trajectories, which are contained in long horizon manipulation tasks. The annotation process involves two primary steps. First, we use Gemini [92] to segment each video based on the sequence of operations and generate detailed text descriptions for each segment. These descriptions accurately capture the operational steps and relevant context. Second, we manually refine Gemini’s annotations regarding the following key aspects:
@@
This thorough process enhances the precision and reliability of the language annotations for the collected trajectories. We show the annotation of a video of a Franka Emika Panda arm picking the apple and placing it in the drawer using the above standard procedure in Figure 5. The results show that our annotation scheme can accurately segment the key actions in the video and provide precise language descriptions of these key actions. More detailed examples of our annotation are provided in the supplementary materials.
@@ IV-A Quantitative Analysis
By breaking down the language descriptions into fine-grained tasks based on verb-noun combinations, RoboMIND includes 479 distinct tasks. In summary, RoboMIND encompasses a range of skills beyond basic manipulations, significantly enhancing the policy model’s manipulation robustness in handling complex and long-horizon tasks.
@@ V-B Single-task Imitation Learning Models
Experimental Results. Figure 12 presents the performance of ACT [112], Diffusion Policy [14], and BAKU [36] across 45 tasks using four types of robots, evaluated in terms of the success rate. In Figure 12, we found that ACT achieves an average success rate of 55.3% across 15 tasks on AgileX, outperforming Franka (30.7%), UR5e (38.0%), and Tien Kung (34.0%). Additionally, ACT also showed promising results on several humanoid robot tasks, including a 60% success rate on HR-CloseDrawerLowerCabinet. These results not only illustrated that ACT shows robust performance in complex dexterous hand manipulation tasks but also underscored the high quality of data gathered in RoboMIND. Similarly, Diffusion Policy also demonstrated its capacity to learn complex tasks, outperforming ACT in several tasks on Franka and Tien Kung. Therefore, we believe that the single-arm, dual-arm, and dexterous hand datasets in RoboMIND can serve as high-quality training sets to improve the performance of single-task imitation learning, thereby advancing the development of the entire imitation learning field. On the other hand, BAKU exhibits lower success rates across most tasks. This discrepancy could be attributed to the hyper-parameter settings from the original BAKU paper, which is primarily optimized for simulation environments rather than real-world robotic platforms. The significant gap between simulation and real-world environments underscores the challenges in directly transferring models from simulated settings to physical robots.
@@ V-C Vision-Language-Action Large Models
Training and Evaluation Setup. We evaluated the performance of three models (OpenVLA [48], RDT-1B [61], and CrossFormer [21]) fine-tuned by the demonstrations from RoboMIND in completing various real-world tasks. Given that the VLA large model exhibits excellent generalization performance, we employed an aggregated dataset sourced from multitask demonstrations for fine-tuning the VLA models. Specifically, we took the official pre-trained VLA models and finetuned them on the multitask datasets for each type of robot, and evaluated their performance on each individual task to determine the extent of generalization achieved, by conducting ten trials for each task. We tested ten trials for each experiment. For OpenVLA [48], which involves fine-tuning the Llama 2 model [94] using a large robotic dataset and adapting it to be a 7-DoF VLA model, we only tested it on the Franka single-arm robot. For RDT-1B and CrossFormer, we tested them on the three types robots.
Experimental Results. Table III presents the success rates for various robot tasks performed using the three different VLA models. The experimental results show that the VLA large models fine-tuned on expert demonstrations from RoboMIND performed well across various different robot tasks. The fine-tuned RDT-1B, compared to CrossFormer and OpenVLA, demonstrated significantly enhanced performance in executing tasks across a range of robot models. This improvement is especially notable for dual-arm manipulation tasks, where RDT-1B excelled. Although the performance of OpenVLA being inferior to that of RDT-1B, it nonetheless achieved a comparable task success rate for straightforward tasks like FR-PlaceBreadPlate and FR-SlideCloseDrawer. CrossFormer, after being fine-tuned with RoboMIND, demonstrated performance improvements in tasks executed by single-arm and humanoid robots.
V-D Generalization of VLA Large Models
Evaluation Setup. We conducted tests to validate the generalization of using RoboMIND to fine-tune the VLA large models, assessing their ability to generalize across real task scenarios with varying backgrounds and different objects of manipulation. Specifically, we evaluated the generalization performance on the FR-PlaceBreadPlate task of OpenVLA [48], RDT-1B [61], and CrossFormer [21] funetuned on the Franka multitask dataset in Section V-C. As shown in Figure 14, we executed the FR-PlaceBreadPlate task on three tablecloths with different unseen background patterns and replaced the grasped bread object with an apple, a banana, and a corn. We tested ten trials for each experiment.
TABLE IV: Generalization results of VLA large models on the FR-PlaceBreadPlate-related tasks.
FR-PlaceBreadPlate 4/10 9/10 10/10
FR-PlaceCornPlate 1/10 5/10 6/10
FR-PlaceBananaPlate 1/10 6/10 9/10
FR-PlaceApplePlate 0/10 3/10 2/10
FR-PlaceBreadPlate (Unseen Background 1) 0/10 1/10 2/10
FR-PlaceBreadPlate (Unseen Background 2) 0/10 1/10 0/10
FR-PlaceBreadPlate (Unseen Background 3) 0/10 0/10 0/10
Experimental Results. As presented in Table IV, both RDT-1B and CrossFormer exhibited good generalizations for manipulating objects, especially for objects like bananas that are similar in shape to the bread-like objects in the training data. However, when it comes to generalizing across unseen backgrounds, RDT-1B, OpenVLA, and CrossFormer performed relatively poorly in the FR-PlaceBreadPlate task.
TABLE V: Success rates of the VLA models before and after training with RoboMIND. The notation ‘(origin)’ indicates models fine-tuned directly on the expert multitask dataset without training on RoboMIND, while ‘(RoboMIND)’ denotes models first trained on the entire RoboMIND dataset and subsequently fine-tuned on the expert multitask dataset.
@@ V-D Generalization of VLA Large Models
V-E Leveraging RoboMIND to Enhance VLA Large Models
@@
Training Setup. Currently, most VLA large models are trained with datasets from robots with arms and grippers and can only be applied to the same types of robots. It is noting that RoboMIND contains valuable data from the Tien Kung humanoid robots with dexterous hands, and we applied this dataset in the pre-training of the RDT-1B and CrossFormer models to enhance their ability to handle real-world tasks that require dexterous hand manipulation by humanoid robots. After that, similar to what we did in Section V-C, we fine-tuned the VLA large models using the expert multitask datasets. Simultaneously, we evaluated whether incorporating RoboMIND into the training would enhance the performance of RDT-1B and CrossFormer on manipulation tasks. We conducted ten tests for each model on each task.
@@
Experimental Results. Table V presents the experimental results of RDT-1B and CrossFormer that were first trained on the entire RoboMIND dataset and then fine-tuned on the expert multitask dataset, compared to those fine-tuned directly on the expert multitask dataset. The results show that training different VLA models using the full RoboMIND dataset led to significant improvements in task success rate across a variety of robot tasks. Especially for dual-arm tasks on the CrossFormer, training with the full RoboMIND dataset significantly enhanced its performance. The training effect improved from being unable to complete each dual-arm task to achieving nearly every test success on AX-TakePotato, AX-PutPepper, and AX-AppleBluePlate. For HR-PressDownToaster from humanoid manipulation tasks, training CrossFormer using RoboMIND also achieved a 100% task success rate. This improvement underscores the robustness and versatility of RoboMIND in facilitating more effective and reliable robotic manipulations.
@@ V-F Failure Case Analysis on Real-world Experiments
V-G Co-training with Real and Simulation Data
To validate the effectiveness of simulation data in RoboMIND, we conducted experiments combining both real-world and simulation data for training. We selected a complex Franka robotic arm task, FR-UprightBlueCup, which requires the robotic arm to rotate nearly 90 degrees, insert its gripper horizontally into the cup’s opening, and restore an overturned cup to its upright position. As shown in Figure 16, we constructed a digital twin simulation environment that closely mirrors the real-world setup, including the robotic arm, table surface, objects, and cameras. We collected 100 real-world trajectories and 500 trajectories in the simulation. We then trained and evaluated the ACT model using different ratios of real-world to simulation data, including real-world data only, simulation data only, and mixed ratios of 100:100, 100:200, 100:300, 100:400, and 100:500. Notably, we did not employ any sim2real transfer techniques but instead directly combined both types of data for co-training.
Figure 17 shows the success rates of ACT in both real-world and simulation environments under different experimental settings. Our observations revealed that increasing the proportion of simulation data improved success rates in both real-world and simulation environments, thanks to our highly accurate simulation environment that closely resembles real-world conditions. However, we also discovered that simulation data alone is insufficient for real-world performance, with real-world data playing a crucial role. For instance, while the combination of 100 real-world trajectories and 500 simulation trajectories achieved a 90% success rate in the simulated environments, using simulation data alone resulted in a dramatic decrease to a 10% success rate in the real world. The primary cause of failure was the cup slipping from the gripper during rotation due to insufficient grip closure. This suggests that significant disparities exist between simulated and real-world physics, particularly for contact-rich tasks, indicating room for improvement in simulation fidelity.

### v3

@@ I Introduction
Beyond establishing such a large-scale and diverse dataset, we conduct extensive experiments to not only validate the dataset’s effectiveness but also evaluate various algorithms’ performance, providing a comprehensive benchmark analysis. Specifically, we evaluate the task success rates using single-task imitation learning methods, including ACT [116], Diffusion Policy [17], and BAKU [39]. Additionally, we assess the generalization capabilities and task success rates of Vision-Language-Action (VLA) large models such as OpenVLA [77], RDT-1B [65], and CrossFormer [24]. The experimental results demonstrate that RoboMIND can be effectively utilized by various single-task imitation learning algorithms and successfully adapted to VLA large models. The high-quality information provided by our dataset enables successful task execution across different approaches in real-world scenarios. Furthermore, pre-training the entire VLA models using the full RoboMIND dataset results in significant improvements in task performance across multiple robot types. To simplify the use of RoboMIND, we provide the code scripts that adapt RoboMIND files with the open-source LeRobot framework [11] at https://github.com/x-humanoid-robomind/x-humanoid-training-toolchain/.
@@ III-A Data Collection and Storage
Teleoperation is widely applied in the data collection processes for various types of robots [85, 114, 108, 40, 63, 106, 84, 16, 100]. Different types of robots also have specific teleoperation devices for data collection. For example, researchers typically use VR headsets and motion capture suits to collect humanoid robot motion data. They capture the state of human movements and map this motion onto the humanoid robot platform, enabling the robot to replicate these movements while simultaneously collecting a comprehensive dataset [16, 100]. RoboMIND contains teleoperation data in real-world and simulation environments from various types of robots, such as single-arm robots (Franka Emika Panda [34], UR5e [88]), dual-arm robots (AgileX Cobot Magic V2.0 [87]), and humanoid robots (X-Humanoid Tien Kung [12]).
@@
Figure 4: We define 8 quality assurance criteria in the data collection process. Touch Excess: Unnecessary contact with objects by the robotic arm; Movement not Smooth: Noticeable jerking or interruptions in robotic arm movements; Secondary Grabbing: Repeated grasping attempts after failures in robotic arm operations; Mechanical Arm Shaking: Abnormal vibrations in the robotic arm; Collision before Grabbing: Collision of the gripper with surrounding objects before grasping; Image Distortion: Data collection quality issues; Failed Placement: Incorrect placement of objects; Gripper out of the Camera: Frames in which the gripper exceeds video frame boundaries. During the data inspection process, all failures were annotated from videos. We show 8 trajectory examples that failed to pass the quality assurance due to different reasons. Each example includes three images that depict the dynamic process of the trajectory. We use red boxes and markers to highlight the reasons for failure.
@@ III-C Data Annotation
While the visual and robot proprioceptive information can be extracted directly from the collected videos and trajectories, we need to provide better semantic information from the data to aid model training. For each collection task, its detailed and accurate linguistic descriptions are provided. These linguistic annotations can be utilized for training currently popular VLA models. In addition, RoboMIND collection tasks encompass numerous long horizon tasks, where a uniform linguistic description may be insufficient to capture the full complexity and nuances of the entire task. Thus, we offer detailed fine-grained linguistic annotations for each movement occurring within a trajectory, as illustrated in Figure 5. We annotate 10k successful robot motion trajectories, which are contained in long horizon manipulation tasks. The annotation process involves two primary steps. First, we use Gemini [96] to segment each video based on the sequence of operations and generate detailed text descriptions for each segment. These descriptions accurately capture the operational steps and relevant context. Second, we manually refine Gemini’s annotations regarding the following key aspects:
@@
This thorough process enhances the precision and reliability of the language annotations for the collected trajectories. We show the annotation of a video of a Franka Emika Panda arm picking the apple and placing it in the drawer using the above standard procedure in Figure 5. The results show that our annotation scheme can accurately segment the key actions in the video and provide precise language descriptions of these key actions. These detailed descriptions can be used for training models like RT-H [7].
@@ IV-A Quantitative Analysis
By breaking down the language descriptions into fine-grained tasks based on verb-noun combinations, RoboMIND includes 479 distinct tasks. In summary, RoboMIND encompasses a range of skills (i.e., verbs from descriptions) beyond basic manipulations, significantly enhancing the policy model’s manipulation robustness in handling complex and long-horizon tasks.
@@ IV-B Qualitative Analysis
The failure data is intended to advance research in areas like failure detection and recovery, data augmentation, and reward generation for reinforcement learning. For instance, training a binary classifier on success/failure data can aid failure detection, SSDF [104] filters high-quality data segments from failure data, and the RLHF approach [6] uses failure data as negative examples to learn accurate patterns. Our failure data can be used seamlessly in these works.
@@ V-B Single-task Imitation Learning Models
Experimental Results. Figure 12 presents the performance of ACT [116], Diffusion Policy [17], and BAKU [39] across 45 tasks using four types of robots, evaluated in terms of the success rate. In Figure 12, we found that ACT achieves an average success rate of 55.3% across 15 tasks on AgileX, outperforming Franka (30.7%), UR5e (38.0%), and Tien Kung (34.0%). Additionally, ACT also showed promising results on several humanoid robot tasks, including a 60% success rate on HR-CloseDrawerLowerCabinet. These results not only illustrated that ACT shows robust performance in complex dexterous hand manipulation tasks but also underscored the high quality of data gathered in RoboMIND. Similarly, Diffusion Policy also demonstrated its capacity to learn complex tasks, outperforming ACT in several tasks on Franka and Tien Kung. Therefore, we believe that the single-arm, dual-arm, and dexterous hand datasets in RoboMIND can serve as high-quality training sets to improve the performance of single-task imitation learning, thereby advancing the development of the entire imitation learning field. On the other hand, BAKU exhibits lower success rates across most tasks. This discrepancy could be attributed to the hyper-parameter settings from the original BAKU paper, which is primarily optimized for simulation environments rather than real-world robotic platforms tested in our experiments. The significant performance gap underscores the challenges in directly transferring models from simulated settings to physical robots.
@@ V-C Vision-Language-Action Large Models
Training and Evaluation Setup. We evaluated the performance of three models (OpenVLA [51], RDT-1B [65], and CrossFormer [24]) fine-tuned by the demonstrations from RoboMIND in completing various real-world tasks. Given that the VLA large model exhibits excellent generalization performance, we employed an aggregated dataset sourced from multitask demonstrations for fine-tuning the VLA models. Specifically, we took the official pre-trained VLA models and fine-tuned them on the multitask datasets for each type of robot, and evaluated their performance on each individual task to determine the extent of generalization achieved, by conducting ten trials for each task. We tested ten trials for each experiment. For OpenVLA [51], which involves fine-tuning the Llama 2 model [98] using a large robotic dataset and adapting it to be a 7-DoF VLA model, we only tested it on the Franka single-arm robot, since the output of OpenVLA is the condition of one end effector and only supports single-arm manipulations.
@@
Experimental Results. Table III presents the success rates for various robot tasks performed using the three different VLA models. The experimental results show that the VLA large models fine-tuned on expert demonstrations from RoboMIND performed well across various different robot tasks. The fine-tuned RDT-1B, compared to CrossFormer and OpenVLA, demonstrated significantly enhanced performance in executing tasks across a range of robot models. This improvement is especially notable for dual-arm manipulation tasks, where RDT-1B excelled. Although the performance of OpenVLA being inferior to that of RDT-1B, it nonetheless achieved a comparable task success rate for straightforward tasks like FR-PlaceBreadPlate and FR-SlideCloseDrawer. CrossFormer, after being fine-tuned with RoboMIND, demonstrated performance improvements in tasks executed by single-arm and humanoid robots, compared with no success in all tasks without fine-tuning.
TABLE IV: Success rates of the VLA models before and after training with RoboMIND. The notation ‘(origin)’ indicates models fine-tuned directly on the expert multitask dataset without training on RoboMIND, while ‘(RoboMIND)’ denotes models first trained on the entire RoboMIND dataset and subsequently fine-tuned on the expert multitask dataset.
@@
V-D Leveraging RoboMIND to Enhance VLA Large Models
@@
Training and Evaluation Setup. Currently, most VLA large models are trained with datasets from robots with arms and grippers and can only be applied to the same types of robots. It is noting that RoboMIND contains valuable data from diverse robots including the Tien Kung humanoid robots with dexterous hands, and we applied this dataset in the pre-training of the RDT-1B and CrossFormer models to enhance their ability to handle real-world dexterous manipulation tasks. After that, we fine-tuned the VLA large models using the expert multitask datasets, which is a small subset (about 1%) of RoboMIND. We conducted ten tests for each model on each task, and we compared the results with those from the official pre-trained and expert data fine-tined models.
@@
Experimental Results. Table IV presents the experimental results of RDT-1B and CrossFormer that were first trained on the entire RoboMIND dataset and then fine-tuned on the expert multitask dataset, compared to those fine-tuned directly on the expert multitask dataset. The results show that training different VLA models using the full RoboMIND dataset led to significant improvements in task success rate across a variety of robot tasks. Especially for dual-arm tasks on the CrossFormer, training with the full RoboMIND dataset significantly enhanced its performance. The training effect improved from being unable to complete each dual-arm task to achieving nearly every test success on AX-TakePotato, AX-PutPepper, and AX-AppleBluePlate. For HR-PressDownToaster from humanoid manipulation tasks, training CrossFormer using RoboMIND also achieved a 100% task success rate. This improvement underscores the robustness and versatility of RoboMIND in facilitating more effective and reliable robotic manipulations.
TABLE V: Success rates of RDT-1B pre-trained with and without the humanoid data.
Success Rate RDT-1B (w/o Humanoid) RDT-1B (w/ Humanoid)
FR-PlaceBreadPlate 9/10 9/10
FR-PickStrawberryInBowl 5/10 6/10 (1/10 $\uparrow$)
FR-OpenTrashCan 5/10 6/10 (1/10 $\uparrow$)
FR-SideCloseDrawer 8/10 8/10
FR-SideOpenDrawer 3/10 5/10 (2/10 $\uparrow$)
Ablation Studies on Humanoid Data. The standardized data across diverse combinations of robotic platforms eases investigations on cross-embodiment generalization and improvements. For example, it is worth exploring whether the humanoid data is helpful for policy learning for other robots, such as the single-arm Franka. We conducted an ablation study by pre-training the RDT-1B model using an incomplete RoboMIND dataset excluding the humanoids data (19k of 107k in RoboMIND) and fine-tuning the model on the expert multitask datasets. Table V shows that the humanoid data enhances the performance of RDT-1B on the unimanual tasks, especially the difficult tasks. Overall, training with the full RoboMIND dataset obtained a 13.3% relative improvement (0.68 v.s. 0.6) of success rates against the incomplete dataset without humanoid data.
V-E Generalization of VLA Large Models
Evaluation Setup. We conducted tests to validate the generalization of using RoboMIND to fine-tune the VLA large models, assessing their ability to generalize across real task scenarios with varying backgrounds and different objects of manipulation. Specifically, we evaluated the generalization performance on the FR-PlaceBreadPlate task of OpenVLA [51], RDT-1B [65], and CrossFormer [24] funetuned on the Franka multitask dataset in Section V-C. Both RDT-1B and CrossFormer are trained on the entire RoboMIND dataset and subsequently fine-tuned using the Franka expert multitask dataset. OpenVLA, in contrast, is directly fine-tuned using the Franka expert multitask dataset. As shown in Figure 14, we executed the FR-PlaceBreadPlate task on three tablecloths with different unseen background patterns and replaced the grasped bread object with an apple, a banana, and a corn. We tested ten trials for each experiment.
TABLE VI: Generalization results of VLA large models on the FR-PlaceBreadPlate-related tasks.
FR-PlaceBreadPlate 4/10 9/10 10/10
FR-PlaceCornPlate 1/10 5/10 6/10
FR-PlaceBananaPlate 1/10 6/10 9/10
FR-PlaceApplePlate 0/10 3/10 2/10
FR-PlaceBreadPlate (Unseen Background 1) 0/10 1/10 2/10
FR-PlaceBreadPlate (Unseen Background 2) 0/10 1/10 0/10
FR-PlaceBreadPlate (Unseen Background 3) 0/10 0/10 0/10
Experimental Results. As presented in Table VI, both RDT-1B and CrossFormer exhibited good generalizations for manipulating objects, especially for objects like bananas that are similar in shape to the bread-like objects in the training data. However, when it comes to generalizing across unseen backgrounds, RDT-1B, OpenVLA, and CrossFormer performed relatively poorly in the FR-PlaceBreadPlate task.
@@ V-F Failure Case Analysis on Real-world Experiments
V-G Real and Simulation Data
To validate the effectiveness of simulation data in RoboMIND, we conducted two sets of experiments.
Co-training with Real and Simulation Data. Firstly, we combined both real-world and simulation data for training. We selected a complex Franka robotic arm task, FR-UprightBlueCup, which requires the robotic arm to rotate nearly 90 degrees, insert its gripper horizontally into the cup’s opening, and restore an overturned cup to its upright position. As shown in Figure 16, we constructed a digital twin simulation environment that closely mirrors the real-world setup, including the robotic arm, table surface, objects, and cameras. We collected 100 real-world trajectories and 500 trajectories in the simulation. We then trained and evaluated the ACT model using different ratios of real-world to simulation data, including real-world data only, simulation data only, and mixed ratios of 100:100, 100:200, 100:300, 100:400, and 100:500. Notably, we did not employ any sim2real transfer techniques but instead directly combined both types of data for co-training. Figure 17 shows the success rates of ACT in both real-world and simulation environments under different experimental settings. Our observations revealed that increasing the proportion of simulation data improved success rates in both real-world and simulation environments, thanks to our highly accurate simulation environment that closely resembles real-world conditions. However, we also discovered that simulation data alone is insufficient for real-world performance, with real-world data playing a crucial role. For instance, while the combination of 100 real-world trajectories and 500 simulation trajectories achieved a 90% success rate in the simulated environments, using simulation data alone resulted in a dramatic decrease to a 10% success rate in the real world. The primary cause of failure was the cup slipping from the gripper during rotation due to insufficient grip closure. This suggests that significant disparities exist between simulated and real-world physics, particularly for contact-rich tasks, indicating room for improvement in simulation fidelity.
@@ V-G Real and Simulation Data
Performance Correlations between Real and Simulation Environments. We further experimented to validate the correlations between performance in the real and simulation environments of the same model We selected five tasks, trained single-task ACT and Diffusion Policy with real-world data, and evaluated them in both the real world and the digital twin simulation environment. Table VII shows positive correlations of the sim-real test results. This aligns with findings in prior research [59] and confirms the positive correlation between simulation and real-world results.
TABLE VII: Success rates of ACT and Diffusion Policy tested in real and simulation environments.
Tests in Different Environments ACT Diffusion Policy
Sim Real Sim Real
FR-PickStrawberryInBowl 3/10 9/10 0/10 5/10
FR-PlaceBreadPlate 2/10 7/10 3/10 8/10
FR-SlideCloseDrawer 2/10 4/10 7/10 10/10
FR-PlacePearBowl 0/10 4/10 4/10 10/10
FR-PlaceBlockPlate 0/10 3/10 2/10 7/10
Pearson Correlation Coefficient 0.83 0.91